When two right angle vectors of magnitude 25 units and 25√3 units find the resultant magnitude

Grade:11

2 Answers

Arun
6 years ago
Dear Ashok
 
Magnitude of Two perpendicular vectors = 25 units  and 25\sqrt3 units
 
Now
Resultant can be solved by using Pythagoras theorem
 
Hence Resultant = \sqrt(625 +1875) = \sqrt2500
= 50 units
 
Regards
Arun (askIITians forum expert)
Sanju
6 years ago
Let a = 25 units and b = 25√3 
If R is the resultant magnitude, the equation for R = [a+ b2]½ 
Substituting the values,
R = [(25)+ (25√3)2]½   = [625+1875]½  = [2500]½  = 50 units
So the required resultant magnitude is 50 units